OpenStudy (mathmath333):

what will be the value of x^1/2.x^1/4.x^1/8... to infinity

OpenStudy (mathmath333):

\[x^\frac{ 1 }{ 2 },x ^\frac{ 1}{ 4 },x^\frac{ 1 }{ 8 }.......\]

OpenStudy (mathmath333):

a.x^2 b.x c.x^3/2 d.x^3 e.none of these

ganeshie8 (ganeshie8):

whats the exact question ?

ganeshie8 (ganeshie8):

are you adding the terms ?

OpenStudy (mathmath333):

here the series could be in AP,GP or HP .....sum of infinite series is the question i think

ganeshie8 (ganeshie8):

wolfram says the sum does not converge http://www.wolframalpha.com/input/?i=%5Csum%5Climits_%7Bk%3D1%7D%5E%7B%5Cinfty%7D+%28x%5E%281%2F2%5Ek%29%29

OpenStudy (mathmath333):

book shows ans as\[x\]

ganeshie8 (ganeshie8):

do you have solution ?

OpenStudy (mathmath333):

no, no hint no solution, therefore i post here

ganeshie8 (ganeshie8):

consider x = 1

ganeshie8 (ganeshie8):

is 1 + 1 + 1 + ... = 1 ?

OpenStudy (mathmath333):

lol

ganeshie8 (ganeshie8):

question is wrong, if possible take a snapshot and attach

OpenStudy (mathmath333):

ok wait

ganeshie8 (ganeshie8):

for example, when x = 3 the sum of first 3 terms itself give u answer very much greater than x http://www.wolframalpha.com/input/?i=%5Csum%5Climits_%7Bk%3D1%7D%5E%7B3%7D+%283%5E%281%2F2%5Ek%29%29

ganeshie8 (ganeshie8):

its not a series, its a product man !

OpenStudy (anonymous):

\[\large \sum_{n = 1}^{\infty}x^{2^{-n}}\]

OpenStudy (anonymous):

He he he.. :)

ganeshie8 (ganeshie8):

\[\huge x^{ \frac{1}{2} +\frac{1}{4} +\frac{1}{8} +\cdots }\]

OpenStudy (anonymous):

This looks nice..

OpenStudy (mathmath333):

oh ,the author was ambigous

OpenStudy (anonymous):

It is a GP with first term 1/2 and common ratio 1/2

ganeshie8 (ganeshie8):

author was clear, you made it ambiguous by replacing \(\huge \cdot \) by a \(\large ,\)

OpenStudy (mathmath333):

ohh !!!

ganeshie8 (ganeshie8):

work out the sum in exponent

OpenStudy (anonymous):

That dot (.) represents Product.. :)

OpenStudy (mathmath333):

the previous question all had (,)..... so i thought ,lol

ganeshie8 (ganeshie8):

stating the problem as clearly as possible and writing your thoughts on the problem really helps

ganeshie8 (ganeshie8):

can you work the sum in exponent?

OpenStudy (mathmath333):

but thanks for that comma ,i have to be cautious ....lol

ganeshie8 (ganeshie8):

:)